DESIGN AN EXCESS 3 TO BCD CODE CONVERTER THAT GIVES OUTPUT CODE 0000 FOR ALL INVALID INPUT...

design an excess 3 to BCD code converter that gives output code 0000 for all invalid input combinations?
Answer #1

A

B

C

D

W

X

Y

Z

0

0

0

0

0

0

0

0

0

0

0

1

0

0

0

0

0

0

1

0

0

0

0

0

0

0

1

1

0

0

0

0

0

1

0

0

0

0

0

1

0

1

0

1

0

0

1

0

0

1

1

0

0

0

1

1

0

1

1

1

0

1

0

0

1

0

0

0

0

1

0

1

1

0

0

1

0

1

1

0

1

0

1

0

0

1

1

1

1

0

1

1

1

0

0

0

1

1

0

0

1

0

0

1

1

1

0

1

0

0

0

0

1

1

1

0

0

0

0

0

1

1

1

1

0

0

0

0

From – k – MAP:

$$ \begin{array}{l} W=\sum m(11,12) \\ =A \bar{B} C D+A B \bar{C} \bar{D} \\ X=\sum m(7,8,9,10) \\ =\bar{A} B C D+A \bar{B} \bar{C}+A \bar{B} \bar{D} \\ Y=\sum m(5,6,9,10) \\ =\bar{A} B \bar{C} D+\bar{A} B C \bar{D}+A \bar{B} \bar{C} D+A \bar{B} C \bar{D} \\ Z=\sum m(4,6,8,10,12) \\ =\bar{A} B \bar{D}+A \bar{B} \bar{D}+A \bar{C} \bar{D} \end{array} $$